# Splitting of uniform bundles on quadrics

### v1

@@ 3.2. The case $l$ is even
We firstly aim to show that $l$ is not even and we will prove it by contradiction. Now suppose that $l$ is even. If $l$ is smaller than $22n-$ 2 and $(l,2n-1)$ is not $(2,5)$, the morphism $\psi_{x}$ (for the definition of $\psi_{x}$, see (3.2)) is constant for each $x\in X$ according to Proposition 2.1. Then $E$ splits. We exclude the remaining cases $l=2n-2$ and $(l,2n-1)=(2,5)$ by calculations.
@@
Suppose $E$ is unsplit, by the same calculation as in [3, Theorem 4.3, Case I], we have $c_{1}(E_{2})=(X_{1}+X_{2})-X_{1}=X_{2}$. Let $f(X_{1},X_{2})$ be a homogeneous polynomial of degree $12n-$ 1 which is symmetrical in $X_{1}$ and $X_{2}$ and represents $c_{2n-1}(E_{1})$. By comparing the coefficients of $T^{0}$ on the left and right sides of the equation (*3), we get
@@
Then we must have $e_{2n}=a$ and $f(X_{1},X_{2})=-a(X_{1}^{2n-2}X_{2}+X_{1}^{2n-4}X_{2}^{3}+\cdots+X_{2}^{2n-1})$, contradicting to the assumption that $f$ is symmetrical in $X_{1}$ and $X_{2}$. ∎
@@
Suppose $E$ is unsplit. Then $a$ in the equation (*3) is not 00 and $\frac{1}{a}E(T,X_{1})=\frac{1}{a}\sum_{k=0}^{6}e_{k}X_{1}^{k}T^{6-k}$ is an approximate solution which has a symmetrical divisor of degree $3$. By Lemma 3.3, we have $\frac{1}{a}e_{6}=0$ or $\frac{1}{a}e_{6}=1$. Let $f(X_{1},X_{2})$ be a homogeneous polynomial representing $c_{3}(E_{1})$.
If $\frac{1}{a}e_{6}$ is 00, we have $f(X_{1},X_{2})\mid R(X_{1},X_{2})(=X_{1}^{6}+X_{1}^{4}X_{2}^{2}+X_{1}^{2}X_{2}^{4}+X_{2}^{6})$. Since the prime factorization of $X_{1}^{6}+X_{1}^{4}X_{2}^{2}+X_{1}^{2}X_{2}^{4}+X_{2}^{6}$ over $\mathbb{Q}[X_{1},X_{2}]$ is $(X_{1}^{4}+X_{2}^{4})(X_{1}^{2}+X_{2}^{2})$, $R(X_{1},X_{2})$ has no divisor symmetrical in $X_{1}$ and $X_{2}$ of degree $3$.
If $\frac{1}{a}e_{6}$ is $1$, then $f(X_{1},X_{2})$ divides $X_{1}^{6}-R(X_{1},X_{2})(=-X_{2}^{2}(X_{1}^{4}+X_{1}^{2}X_{2}^{2}+X_{2}^{4}))$. The prime factorization of $X_{1}^{4}+X_{1}^{2}X_{2}^{2}+X_{2}^{4}$ over $\mathbb{Q}[X_{1},X_{2}]$ is $(X_{1}^{2}+X_{1}X_{2}+X_{2}^{2})(X_{1}^{2}-X_{1}X_{2}+X_{2}^{2})$. Therefore, $X_{1}^{6}-R(X_{1},X_{2})$ has no divisor symmetrical in $X_{1}$ and $X_{2}$ of degree $3$.
@@ 3.3. The case $l$ is odd
By [9, Proposition 2.1], $c_{l+1}(\psi_{x}^{*}U^{\vee})$ and $c_{2n-1-l}(\psi_{x}^{*}Q)$ are numerically non-negative, hence both $a_{l+1}$ and $b_{2n-1-l}$ are non-negative. As $-a$ is $a_{l+1}b_{2n-1-l}$ and $a$ is not 00, we have $-a>0$. So for any $t\in\mathbb{R}$, $t^{2n}-a$ is bigger than 00. In other words, $S_{1}(t,0,1)$ and $S_{2}(t,0,1)$ are non-zero for any $t\in\mathbb{R}$. By the equations $E(T,1)-aR(1,0)=S_{1}(T,1,0)S_{2}(T+1,1,0)=S_{1}(T,0,1)S_{2}(T+1,0,1)$, for any $t\in\mathbb{R}$, $E(t,1)-aR(1,0)$ is not 00. ∎
@@
Recall the equation $\frac{1}{a}E(t,1)-R(1,z)=\frac{1}{a}S_{1}(t,1,z)S_{2}(t+1,1,z)$. Note that $l$ is odd implies $l+1\geq 2n-1-l>1$. So by Lemma 3.3, we have $\frac{1}{a}e_{2n}=0$ or $\frac{1}{a}e_{2n}=1$. And when $\frac{1}{a}e_{2n}$ is 00, $\frac{1}{a}E(0,1)-R(1,\exp\frac{2i\pi}{2n+2})$ vanishes; when $\frac{1}{a}e_{2n}$ is $1$, $\frac{1}{a}E(0,1)-R(1,\exp\frac{2i\pi}{2n})$ vanishes.
Suppose that $n$ is $3$. When $\frac{1}{a}e_{2n}$ is 00, we have $\frac{1}{a}S_{1}(0,1,z)S_{2}(1,1,z)=-(z^{6}+z^{4}+z^{2}+1)$. The prime factorization of $z^{6}+z^{4}+z^{2}+1$ over $\mathbb{Q}[z]$ is $(z^{4}+1)(z^{2}+1)$. Note $\deg(S_{1})\geq\deg(S_{2})$ and $\deg(S_{2})>1$, we must have $S_{1}(0,1,z)=\lambda(z^{4}+1)$ for some $\lambda\in\mathbb{Q}$. Then $\exp\frac{2i\pi}{8}$ is a root of $S_{1}(0,1,z)$ and hence $\frac{1}{a}E(T,X_{1})$ is primitive. When $\frac{1}{a}e_{2n}$ is $1$, we have $\frac{1}{a}S_{1}(0,1,z)S_{2}(1,1,z)=-(z^{6}+z^{4}+z^{2})=-z^{2}(z^{4}+z^{2}+1)=-z^{2}(z^{2}+z+1)(z^{2}-z+1)$. Note that $S_{1}$ is symmetrical in $X_{1},X_{2}$ and $\deg(S_{1})$ is at least $\deg(S_{2})$, we have $S_{1}(0,1,z)=\lambda(z^{4}+z^{2}+1)$. Then $\exp\frac{2i\pi}{6}$ is a root of $S_{1}(0,1,z)$ and hence $\frac{1}{a}E(T,X_{1})$ is also primitive.
Suppose now $n$ is at least $4$ and that $\frac{1}{a}E(T,X_{1})$ is not a primitive solution. Then we have
$S_{2}(1,1,\exp\frac{2i\pi}{2n+2})=0\text{ or }S_{2}(1,1,\exp\frac{2i\pi}{2n})=0.$ (3.3)
@@
(Note that for the last inequality, we use the condition $n\geq 4$). Since the roots of $S_{2}(0,1,z)$ satisfy $z^{2n}=1$ or $z^{2n+2}=1$, it suffices to show that $S_{2}(0,1,z)$ has a non-zero root $y_{0}$ with argument satisfying $\frac{\pi}{2n}<\arg y_{0}<\frac{3\pi}{2n}$. By the above inequalities, we have $y_{0}=\frac{2i\pi}{2n+2}$ or $y_{0}=\frac{2i\pi}{2n}$. Then $\frac{1}{a}E(T-X_{1},X_{1})$ is a primitive approximate solution.
@@
Now we complete the proof for the case $l$ is odd. By Proposition 3.5 and Proposition 3.9, we have the following possibilities (note the coefficient of $T^{2n}$ in $E(T,X_{1})$ is $1$):
@@
$X_{1}-\exp\frac{2i\pi}{2n+2}X_{2}\mid R(X_{1},X_{2}),~{}(X_{1}-\exp\frac{2i\pi}{2n+2}X_{2})^{2}\nmid R(X_{1},X_{2}),~{}X_{1}-\exp\frac{2i\pi}{2n+2}X_{2}\nmid 1.$
By Eisenstein’s criterion, $T^{2n}-aR(X_{1},X_{2})$ is irreducible as the polynomial in the variable $T$ with coefficients in $\mathbb{Q}[X_{1},X_{2}]$. This leads to a contradiction. Similar arguments can be applied to the case $E(T-X_{1},X_{1})$ is $T^{2n}$.
If $E(T,X_{1})$ is $b_{1}^{n}\Sigma_{n}(\frac{T^{2}}{b_{1}},X_{1}^{2})(=\Sigma_{n}(T^{2},b_{1}X_{1}^{2}))$, then we have the equalities $E(T,X_{1})-aR(X_{1},X_{2})=\Sigma_{n}(T^{2},b_{1}X_{1}^{2})-\Sigma_{n}(b_{1}X_{1}^{2},b_{1}X_{2}^{2})=(T^{2}-b_{1}X_{2}^{2})\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ (for the last equality, see [2, Section 7.2] for example). We will make use of the following claim, whose proof will be given in Proposition A.3.
@@
As $b_{1}$ is not 00, $\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ is irreducible. We have $(T^{2}-b_{1}X_{2}^{2})\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})=S_{1}(T,X_{1},X_{2})S_{2}(T+X_{1},X_{1},X_{2})$. Since $S_{1}(T,X_{1},X_{2})$ is symmetrical in $X_{1},X_{2}$ and the coefficient of $T^{2n}$ of $S_{1}$ is $1$, then $S_{1}(T,X_{1},X_{2})$ is $\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ and hence $S_{2}(T+X_{1},X_{1},X_{2})$ is $T^{2}-b_{1}X_{2}^{2}$. So $S_{2}(T,X_{1},X_{2})$ is $(T-X_{1})^{2}-b_{1}X_{2}^{2},$ which is not symmetrical in $X_{1},X_{2}.$ We get a contradiction.
@@
Every uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ splits.
@@
We have shown that there is no unsplit uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$ for $n\geq 3$. Now let $E^{\prime}$ be a uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+2}$. For every smooth hyperplane section $\mathbb{Q}^{2n+1}\hookrightarrow\mathbb{Q}^{2n+2}$, the restriction $E^{\prime}|_{\mathbb{Q}^{2n+1}}$ splits. Then we conclude by [6, Corollary 3.3]. ∎
Given a generalised Grassmannian $X$, we denote by $\mu(X)$ the maximum positive integer verifying that every uniform vector bundle whose rank is at most $\mu(X)$ splits. And we call $\mu(X)$ the splitting threshold for uniform vector bundles on $X$. For example, $\mu(\mathbb{P}^{n})$ is $n-1$. The main theorem in [3] tells us that for most generalised Grassmannians $X$, $\mu(X)$ is equal to $e.d.(\mathrm{VMRT})$ (see [3, Page 3, Table 2]). Note that for $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}$, their $e.d.(\mathrm{VMRT})$ are all $12n-$ 1.
@@
  1. (1)
    Theorem 3.11 reveals that the splitting thresholds for uniform vector bundles on $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ are at least $2n$. Actually, $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ are the first known examples such that $\mu(X)$ is bigger than $e.d.(\mathrm{VMRT})$.
  2. (2)
    The threshold $\mu(\mathbb{Q}^{2n+1})$ is $2n$, as the tangent bundle $T_{\mathbb{Q}^{2n+1}}$ is unsplit. Moreover, the splitting type of $T_{\mathbb{Q}^{2n+1}}$ is $(2,1,\dots,1,0)$.
In [3], the authors classify the unsplit uniform bundles of minimal ranks on the generalised Grassmannians $B_{n}/P_{k}~{}(2\leq k<\frac{2n}{3})$, $B_{n}/P_{n-2}$, $B_{n}/P_{n-1}$ and $D_{n}/P_{k}~{}(2\leq k<\frac{2n-2}{3})$, $D_{n}/P_{n-3}$, $D_{n}/P_{n-2}$. As direct corollaries of Theorem 3.11, we can give further classification results for uniform bundles on $B_{n}/P_{k}~{}(\frac{2n}{3}\leq k\leq n-3)$ and $D_{n}/P_{k}~{}(\frac{2n-2}{3}\leq k\leq n-4)$.
@@
Let $X$ be $B_{n}/P_{k}$, where $k$ is $\frac{2n}{3}$ and $k$ is at least $6$. Let $E$ be a uniform vector bundle on $X$ of rank $r$.
@@
Note that the $e.d.(\mathrm{VMRT})$ of $X$ is $k-1(=2n-2k-1)$. By [3, Theorem 1.1 (1)], the first assertion follows. For the case $r=k$, since $12n-2k+1=k+$ 1 is at least $7$, $E|_{\mathbb{Q}^{2n-2k+1}}$ splits by Theorem 3.11. Suppose $E$ is unsplit, then $E|_{\mathbb{P}^{k}}$ is also unsplit. Then the second assertion follows from [3, Proposition 4.4]. ∎
@@
  • If $r$ is smallet than $k$, then $E$ is a direct sum of line bundles.

### v2

@@ 3.2. The case $l$ is even
We first show that $l$ is not even and we will prove it by contradiction. Now suppose that $l$ is even. If $l$ is smaller than $22n-$ 2 and $(l,2n-1)$ is not $(2,5)$, the morphism $\psi_{x}$ (for the definition of $\psi_{x}$, see (3.2)) is constant for each $x\in X$ according to Proposition 2.1. Then $E$ splits. We exclude the remaining cases $l=2n-2$ and $(l,2n-1)=(2,5)$ by calculations.
@@
Suppose $E$ is unsplit, by the same calculation as in [4, Theorem 4.3, Case I], we have $c_{1}(E_{2})=(X_{1}+X_{2})-X_{1}=X_{2}$. Let $f(X_{1},X_{2})$ be a homogeneous polynomial of degree $12n-$ 1 which is symmetric in $X_{1}$ and $X_{2}$ and represents $c_{2n-1}(E_{1})$. By comparing the coefficients of $T^{0}$ on the left and right sides of the equation (*3), we get
@@
Then we must have $e_{2n}=a$ and $f(X_{1},X_{2})=-a(X_{1}^{2n-2}X_{2}+X_{1}^{2n-4}X_{2}^{3}+\cdots+X_{2}^{2n-1})$, contradicting the assumption that $f$ is symmetric in $X_{1}$ and $X_{2}$. ∎
@@
Suppose $E$ is unsplit. Then $a$ in the equation (*3) is not 00 and $\frac{1}{a}E(T,X_{1})=\frac{1}{a}\sum_{k=0}^{6}e_{k}X_{1}^{k}T^{6-k}$ is an approximate solution which has a symmetric divisor of degree $3$. By Lemma 3.3, we have $\frac{1}{a}e_{6}=0$ or $\frac{1}{a}e_{6}=1$. Let $f(X_{1},X_{2})$ be a homogeneous polynomial representing $c_{3}(E_{1})$.
If $\frac{1}{a}e_{6}$ is 00, we have $f(X_{1},X_{2})\mid R(X_{1},X_{2})(=X_{1}^{6}+X_{1}^{4}X_{2}^{2}+X_{1}^{2}X_{2}^{4}+X_{2}^{6})$. Since the prime factorization of $X_{1}^{6}+X_{1}^{4}X_{2}^{2}+X_{1}^{2}X_{2}^{4}+X_{2}^{6}$ over $\mathbb{Q}[X_{1},X_{2}]$ is $(X_{1}^{4}+X_{2}^{4})(X_{1}^{2}+X_{2}^{2})$, $R(X_{1},X_{2})$ has no divisor symmetric in $X_{1}$ and $X_{2}$ of degree $3$.
If $\frac{1}{a}e_{6}$ is $1$, then $f(X_{1},X_{2})$ divides $X_{1}^{6}-R(X_{1},X_{2})(=-X_{2}^{2}(X_{1}^{4}+X_{1}^{2}X_{2}^{2}+X_{2}^{4}))$. The prime factorization of $X_{1}^{4}+X_{1}^{2}X_{2}^{2}+X_{2}^{4}$ over $\mathbb{Q}[X_{1},X_{2}]$ is $(X_{1}^{2}+X_{1}X_{2}+X_{2}^{2})(X_{1}^{2}-X_{1}X_{2}+X_{2}^{2})$. Therefore, $X_{1}^{6}-R(X_{1},X_{2})$ has no divisor symmetric in $X_{1}$ and $X_{2}$ of degree $3$.
@@ 3.3. The case $l$ is odd
By [13, Proposition 2.1], $c_{l+1}(\psi_{x}^{*}U^{\vee})$ and $c_{2n-1-l}(\psi_{x}^{*}Q)$ are numerically non-negative, hence both $a_{l+1}$ and $b_{2n-1-l}$ are non-negative. As $-a$ is $a_{l+1}b_{2n-1-l}$ and $a$ is not 00, we have $-a>0$. So for any $t\in\mathbb{R}$, $t^{2n}-a$ is bigger than 00. In other words, $S_{1}(t,0,1)$ and $S_{2}(t,0,1)$ are non-zero for any $t\in\mathbb{R}$. By the equations $E(T,1)-aR(1,0)=S_{1}(T,1,0)S_{2}(T+1,1,0)=S_{1}(T,0,1)S_{2}(T+1,0,1)$, for any $t\in\mathbb{R}$, $E(t,1)-aR(1,0)$ is not 00. ∎
@@
Recall the equation $\frac{1}{a}E(t,1)-R(1,z)=\frac{1}{a}S_{1}(t,1,z)S_{2}(t+1,1,z)$. Note that $l$ is odd implies $l+1\geq 2n-1-l>1$. So by Lemma 3.3, we have $\frac{1}{a}e_{2n}=0$ or $\frac{1}{a}e_{2n}=1$. And when $\frac{1}{a}e_{2n}$ is 00, $\frac{1}{a}E(0,1)-R(1,\exp\frac{2\pi i}{2n+2})$ vanishes; when $\frac{1}{a}e_{2n}$ is $1$, $\frac{1}{a}E(0,1)-R(1,\exp\frac{2\pi i}{2n})$ vanishes.
Suppose that $n$ is $3$. When $\frac{1}{a}e_{2n}$ is 00, we have $\frac{1}{a}S_{1}(0,1,z)S_{2}(1,1,z)=-(z^{6}+z^{4}+z^{2}+1)$. The prime factorization of $z^{6}+z^{4}+z^{2}+1$ over $\mathbb{Q}[z]$ is $(z^{4}+1)(z^{2}+1)$. Note $\deg(S_{1})\geq\deg(S_{2})$ and $\deg(S_{2})>1$, we must have $S_{1}(0,1,z)=\lambda(z^{4}+1)$ for some $\lambda\in\mathbb{Q}$. Then $\exp\frac{2\pi i}{8}$ is a root of $S_{1}(0,1,z)$ and hence $\frac{1}{a}E(T,X_{1})$ is primitive. When $\frac{1}{a}e_{2n}$ is $1$, we have $\frac{1}{a}S_{1}(0,1,z)S_{2}(1,1,z)=-(z^{6}+z^{4}+z^{2})=-z^{2}(z^{4}+z^{2}+1)=-z^{2}(z^{2}+z+1)(z^{2}-z+1)$. Note that $S_{1}$ is symmetric in $X_{1},X_{2}$ and $\deg(S_{1})$ is at least $\deg(S_{2})$, we have $S_{1}(0,1,z)=\lambda(z^{4}+z^{2}+1)$. Then $\exp\frac{2\pi i}{6}$ is a root of $S_{1}(0,1,z)$ and hence $\frac{1}{a}E(T,X_{1})$ is also primitive.
Suppose now $n$ is at least $4$ and $\frac{1}{a}E(T,X_{1})$ is not a primitive solution. Then we have
$S_{2}(1,1,\exp\frac{2\pi i}{2n+2})=0\text{ or }S_{2}(1,1,\exp\frac{2\pi i}{2n})=0.$ (3.3)
@@
(Note that for the last inequality, we use the condition $n\geq 4$). Since the roots of $S_{2}(0,1,z)$ satisfy $z^{2n}=1$ or $z^{2n+2}=1$, it suffices to show that $S_{2}(0,1,z)$ has a non-zero root $y_{0}$ with argument satisfying $\frac{\pi}{2n}<\arg y_{0}<\frac{3\pi}{2n}$. By the above inequalities, we have $y_{0}=\frac{2\pi i}{2n+2}$ or $y_{0}=\frac{2\pi i}{2n}$. Then $\frac{1}{a}E(T-X_{1},X_{1})$ is a primitive approximate solution.
@@
Now we complete the proof for the case $l$ is odd. By Proposition 3.5 and Proposition 3.9, we have the following possibilities (note that the coefficient of $T^{2n}$ in $E(T,X_{1})$ is $1$):
@@
$X_{1}-\exp\frac{2\pi i}{2n+2}X_{2}\mid R(X_{1},X_{2}),~{}(X_{1}-\exp\frac{2\pi i}{2n+2}X_{2})^{2}\nmid R(X_{1},X_{2}),~{}X_{1}-\exp\frac{2\pi i}{2n+2}X_{2}\nmid 1.$
By Eisenstein’s criterion, $T^{2n}-aR(X_{1},X_{2})$ is irreducible considered as the polynomial in the variable $T$ with coefficients in $\mathbb{Q}[X_{1},X_{2}]$. This leads to a contradiction. Similar arguments can be applied to the case $E(T-X_{1},X_{1})$ is $T^{2n}$.
If $E(T,X_{1})$ is $b_{1}^{n}\Sigma_{n}(\frac{T^{2}}{b_{1}},X_{1}^{2})(=\Sigma_{n}(T^{2},b_{1}X_{1}^{2}))$, then we have the equalities $E(T,X_{1})-aR(X_{1},X_{2})=\Sigma_{n}(T^{2},b_{1}X_{1}^{2})-\Sigma_{n}(b_{1}X_{1}^{2},b_{1}X_{2}^{2})=(T^{2}-b_{1}X_{2}^{2})\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ (for the last equality, see [2, Section 7.2] for example). We will make use of the following claim, whose proof will be given in Proposition B.3.
@@
As $b_{1}$ is not 00, $\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ is irreducible. We have $(T^{2}-b_{1}X_{2}^{2})\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})=S_{1}(T,X_{1},X_{2})S_{2}(T+X_{1},X_{1},X_{2})$. Since $S_{1}(T,X_{1},X_{2})$ is symmetric in $X_{1},X_{2}$ and the coefficient of $T^{2n}$ of $S_{1}$ is $1$, then $S_{1}(T,X_{1},X_{2})$ is $\Sigma_{n-1}(T^{2},b_{1}X_{1}^{2},b_{1}X_{2}^{2})$ and hence $S_{2}(T+X_{1},X_{1},X_{2})$ is $T^{2}-b_{1}X_{2}^{2}$. So $S_{2}(T,X_{1},X_{2})$ is $(T-X_{1})^{2}-b_{1}X_{2}^{2},$ which is not symmetric in $X_{1},X_{2}.$ We get a contradiction.
@@
Assume $n$ is at least $3$, then every uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$ or $\mathbb{Q}^{2n+2}$ splits and $\mu(\mathbb{Q}^{2n+1})$ is $2n$.
@@
Combining Proposition 2.1, Proposition 3.6, Proposition 3.7 and Proposition 3.10, we can show that every uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$ splits for $n\geq 3$. As the tangent bundle $T_{\mathbb{Q}^{2n+1}}$ is unsplit, the threshold $\mu(\mathbb{Q}^{2n+1})$ is $2n$.
Now let $E^{\prime}$ be a uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+2}$. For every smooth hyperplane section $\mathbb{Q}^{2n+1}\hookrightarrow\mathbb{Q}^{2n+2}$, the restriction $E^{\prime}|_{\mathbb{Q}^{2n+1}}$ is a uniform bundle of rank $2n$ on $\mathbb{Q}^{2n+1}$, hence $E^{\prime}|_{\mathbb{Q}^{2n+1}}$ splits. So by [9, Corollary 3.3], $E^{\prime}$ splits. ∎
@@
For a majority of generalized Grassmannians $X$, $\mu(X)$ is $e.d.(\mathrm{VMRT})$ (see [4, Page 3, Table 2]). Note that the $e.d.(\mathrm{VMRT})$ of both $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}$ are $12n-$ 1. Theorem 3.11 shows that the splitting thresholds for uniform vector bundles on $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ are at least $2n$. So $\mathbb{Q}^{2n+1}$ and $\mathbb{Q}^{2n+2}(n\geq 3)$ are the first known examples such that $\mu(X)$ is bigger than $e.d.(\mathrm{VMRT})$.
In [4], the authors classify all unsplit uniform bundles of minimal rank on the generalized Grassmannians $B_{n}/P_{k}~{}(2\leq k<\frac{2n}{3})$, $B_{n}/P_{n-2}$, $B_{n}/P_{n-1}$ and $D_{n}/P_{k}~{}(2\leq k<\frac{2n-2}{3})$, $D_{n}/P_{n-3}$, $D_{n}/P_{n-2}$. As direct corollaries of Theorem 3.11, we can give further classification results for uniform bundles on $B_{n}/P_{k}~{}(\frac{2n}{3}\leq k\leq n-3)$ and $D_{n}/P_{k}~{}(\frac{2n-2}{3}\leq k\leq n-4)$.
@@
Let $X$ be $B_{n}/P_{k}$, where $k$ is $\frac{2n}{3}$ and is at least $6$. Let $E$ be a uniform vector bundle on $X$ of rank $r$.
@@
Note that $e.d.(\mathrm{VMRT})$ of $X$ is $k-1(=2n-2k-1)$. By [4, Theorem 1.1 (1)], the first assertion follows. For the case $r=k$, since $12n-2k+1=k+$ 1 is at least $7$, $E|_{\mathbb{Q}^{2n-2k+1}}$ splits by Theorem 3.11. Suppose $E$ is unsplit, then $E|_{\mathbb{P}^{k}}$ is also unsplit. The second assertion then follows from [4, Proposition 4.4]. ∎
@@
  • If $r$ is smaller than $k$, then $E$ is a direct sum of line bundles.